Copier a can print 20 color pages per minute copier b can print 18 color pages per minute. If both machines begin printing at the same instant, how long will it be until they again begin printing a page exactely at the same time?
Answer by checkley77(12844)

1/20 & 1/18 find thew LCD
20=2*2*5
18=2*3*3
Thus the LCD=2*5*3*3=90
The 90th page will be printed at the same time.